OpenStudy (dls):

If g is inverse of f and f'(x)=1/1+x^3 then find g'(x).

OpenStudy (dls):

@zepdrix @experimentX

OpenStudy (experimentx):

whadoyou mean?? \[ f^{-1}(x) = \frac{1}{1+x^3}\] and \[ g(x) = f^{-1}(x)\] ??

OpenStudy (dls):

yes correct

OpenStudy (experimentx):

then \[ g(x) = \frac{1}{1+x^3}\] find the derivative ... what's the problem?

OpenStudy (dls):

um sorry its f'(x) not f inverse x

OpenStudy (experimentx):

straight way, this looks bad http://www.wolframalpha.com/input/?i=Integrate+1%2F%281%2Bx%5E3%29

OpenStudy (dls):

we are not supposed to integrate anything here..

zepdrix (zepdrix):

Derivative of the Inverse Function: \[\large \frac{d}{dx}f^{-1}(x)=\frac{1}{f^{-1}\left[f'(x)\right]}\] I think we can use this somehow.. \[\large g^{-1}(x)=f(x)\] \[\large \frac{d}{dx}g^{-1}(x)\qquad=\qquad\frac{d}{dx}f(x)\qquad=\qquad\frac{1}{f^{-1}\left[f'(x)\right]}\] Hmm I'm not doing this correctly... but I think this problem will involve that formula somehow... thinkinggg

zepdrix (zepdrix):

Oh i wrote that incorrectly,\[\large \frac{d}{dx}g(x)=\frac{d}{dx}f^{-1}(x)=...\]Yah i think this will work.

zepdrix (zepdrix):

\[\large =\frac{1}{f^{-1}\left[f'(x)\right]}\]

OpenStudy (dls):

How about this?? Multiplying f(x) on both sides.. \[\large f(g(x))=x\] Differentiating.. \[\Large f'(g(x)).g'(x)=1\] \[\Large g'(x)=1+g(x)^3\]

OpenStudy (experimentx):

is 0 on option?

zepdrix (zepdrix):

\[\large g'(x)=\frac{1}{f^{-1}\left[\dfrac{1}{1+x^3}\right]}\]

OpenStudy (dls):

nope,ans is w hat i wrote

zepdrix (zepdrix):

Hmm

OpenStudy (experimentx):

yes yes ... sorry, i equated with 1

OpenStudy (dls):

thanks! :)